Find the quadratic equation whose roots are -2/3 and -1/4


1
Expert's answer
2022-03-01T15:51:44-0500

Let the quadratic equation be


"ax^2+bx+c=0, a\not=0"

By the Vieta's theorem


"x_1+x_2=-b\/a"

"x_1x_2=c\/a"

Let "a=1." Then


"b=-(-2\/3-1\/4)=11\/12"

"c=(-2\/3)(-1\/4)=1\/6"

The quadratic equation is


"x^2+\dfrac{11}{12}x+\dfrac{1}{6}=0"

Or


"12x^2+11x+2=0"